OpenStudy (loser66):

Is there any link between gamma distribution and normal distribution? Please, help

OpenStudy (ikram002p):

ok , im not expert but in my opinion i think they are different things , not derived from each other although both are used to convert sample to universe .

OpenStudy (loser66):

Thanks friend, I review for final, just want to make the life easier. :)

OpenStudy (ikram002p):

good luck all the best :)

OpenStudy (loser66):

:)

OpenStudy (loser66):

WWWWWWWWWWWHHHHHHHHHAAAt?? I don't want to study anymore. I go to sell Hoagies. :)

OpenStudy (ikram002p):

http://prntscr.com/59nfra

ganeshie8 (ganeshie8):

lol

OpenStudy (loser66):

\[f(x)= \frac{1}{2\sqrt{2\pi}}e^{(-1/8)x^2}\] \[F(x)=\int_{-\infty}^{x} f(t)dt=\int_{-\infty}^{x}\frac{1}{2\sqrt{2\pi}}e^{(-1/8)t^2}dt\] \(W = X^2\) \[G(W)= F(\sqrt w)- F(-\sqrt w)\] \[G'(W)= F'(\sqrt w)- F'(-\sqrt w)\] the last line is \[G'(W) = g(w) = f(\sqrt w) ((1/2)w^{-1/2} - f(-\sqrt w)((1/2)((-1/2)w^{-1/2}\] I don't get the last line, please, explain me

OpenStudy (loser66):

I type the last line wrong, it is

OpenStudy (loser66):

\[G'(W) = g(w) = f(\sqrt w) ((1/2)w^{-1/2}) - f(-\sqrt w)((-1/2)w^{-1/2})\]

OpenStudy (loser66):

I have group study now, appreciate any tips. Please, leave your guide here. I'll be back right after class. :)

OpenStudy (dumbcow):

the last line is derived from using chain rule \[f(x) = F'(x)\] \[(\pm \sqrt{w})' = \pm (1/2)w^{-1/2}\] regarding your question on distributions the gamma distribution can resemble the normal distribution as k-> infinity "For large k the gamma distribution converges to Gaussian distribution with mean μ = kθ and variance σ2 = kθ2." http://en.wikipedia.org/wiki/Gamma_distribution#Others

OpenStudy (loser66):

Thank you. :)